A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100lb of a mixture that will sell for 11.25 per pound. how much of each type of bean should be used in the mixture ? label your answers
Answer by checkley71(8403)

100*11.25=9X+12(100-X) OR 1125=9X+1200-12X OR 3X=75 OR X=25 LBS OF $9 COFFEE
THE WE HAVE 100-25=75 LBS OF $12 COFFEE